Show that a function of bounded variation on [a,b] is bounded therein


1
Expert's answer
2021-03-15T08:10:58-0400

 Let f be a function of bounded variation on the interval [a,b].

Then there exists a positive real number M>0 such that for all partitions P∈ P[a,b]

 we have that:


"V_f(P)=\Sigma_{k=1}^n|f(x_k)-f(x_{k-1})|\leq M"

For all x∈[a,b] consider the partition P={a,x,b} (where P={a,b} if x=a or x=b).

Then:


"V_f(\{a,x,b\})=|f(b)-f(x)|+|f(x)-f(a)|\leq M"

Hence we have that "|f(x)-f(a)|\leq M," so for all x∈[a,b] we have that "|f(x)|\leq |f(a)|+M" ,

So, f is bounded on [a,b]